# What Makes a Visualization Complex?

### v1

@@ IV Comparison to Previous Works
IV.1 Network-based Visual Complexity: Compared to Purchase et al.
@@
To replicate and understand this extended large dataset, and compare results with the Purchase et al. study on network aesthetics, we curated a subset of our dataset containing network-style visualizations (primarily node-link diagrams, as shown in Apdx. Figure 20). To better match the original study design, we manually (1) filtered for direct node-link diagram (such that the matrix views and treemaps were removed; (2) Removed all text and annotations using automated masking followed by a manual cleanup to isolate pure graphical elements given that these texts may activate the edge detector. We refer to this cleaned subset as HP-Trees and Networks. Using our 11 remaining metrics (excluding O.TiR), our PLS regression achieved an of 0.58 indicating a strong fit—substantially higher than the 0.25 reported in Purchase et al. [107]. Notably, O.ED (edge density, ), O.FP (feature points, ), O.MeC (), and O.CF () were key contributors to perceived complexity in this subset (Figure 7). We would credit the results to the addition of O.MeC and O.FP, which emerge as notable contributors to perceived VC in our dataset.
IV.2 Information Theoretic Measure of Color and Texture on Clutter Effect: Compared to Rosenholtz et al.
Rosenholtz et al. designed this clutter-based metric to capture color and texture distributions in a natural scene. Our dataset contains 311 ‘Grid and Matrix’ or ‘heatmaps’ [11], of which 185 are spatial, continuous, where spatial location ‘is given’ [96], making them more closely aligned with Rosenholtz’s definition of structured scenes. The rest are discrete, with pixel locations artificially generated—where structured grids and matrix introduce corners and edges. This contrast is evident in our dataset in these two sets of heatmaps (Figure 8). We also removed all texts before modeling these two sets of visualization images to avoid edges and corners confounded by the presence of texts.
@@
In summary, the continuity-specific results largely mirror the overall findings in terms of dominant contributing metrics, with some variation based on the stimulus type the metric models. What makes this particularly intriguing is that O.FC is highly effective in explaining complexity for continuous heatmap images but had limited explanatory power for artificially generated heatmaps. Given that the coefficients are higher by individual visualization types and continuity, we recommend that type categorization for VC must take into account continuity.
@@ V General Discussion
A central problem for the science of visualization is related to how perceptual experience influences people’ understanding about data. One of such high-level perceptual experience is VC, captured in task-free conditions. Our present experiment contributes to understanding the alignment between human reading and metric measurement. This section discusses theoretical and design implications of our results.
@@ V.1 Comparisons to Natural Scene Understanding in terms of Item-Count and Feature Congestion
Multiscale structures emerge in our study. Number of elments operates on object-associated O.MeC and O.TiR, and low-level features such as edges O.ED and corners O.FP, contributing significantly to VC across overall dataset. This result indicates that human perception does not operate on a single scale [151]. This high-level perception seems to align to Turing’s distinct hypothesis: patterned structures can self-organize through local competitive interactions [136]. As a result, perception can be fast and minimizes the need for generic instruction to perform pattern findings. Turing described that people’s perceptual computing in this competing mechanism appeared to be more robust to noise, where global structural emergence exhibits topological robustness of insensitivity to variations and noises. In visualization contexts, global structures formed by color and pattern may support visual Gestalt principles, as captured by the two object-based metrics: O.TiR (related to text) and O.MeC (related to color and object boundaries).
These size-scale representations are not limited to low-level elements such as points and edges, but also emerge at the level of higher-order textures. This is particularly evident in heatmaps where spatial location is given and aligns with the descriptor attribute O.FC, as demonstrated in our experiments (Figure 8). In this context, O.FC represents a meaningful step forward in quantifying human perceptual experiences by treating visualizations as statistical, scene-based representations.
@@ V.2 Graphic or Text? The Influence of Semantic Information on Perceived Visual Complexity
Our findings suggest that the semantic information conveyed by text in visualizations impacts perceived VC differently than graphic elements. We conjectured that explanatory text may reduce perceived VC by enhancing understandability, e. g., Hearst et al. also highlighted the significance of text, considering it as equally important as the visualizations themselves [49]. We found a positive effect of text on VC to a certain extent. In general, explanatory text reduces perceived VC due to its contribution to understandability. However, when the amount of text becomes excessive, it may increase complexity, shifting from aiding comprehension to overwhelming the visual features.
@@
Analyzing variations in text layout and related attributes may reveal new ways to understand perceived VCs. Object-attribute-level experience has been advocated by vision scientists to the extent that they question the usefulness of low-level features [117]. Previous eye-tracking behavior data for capturing memorability suggested that people can quickly distinguish text and titles. These text elements always occupied separate areas on the screen, making them easy for the human visual system to decode. In previous studies on three-dimensional scenes, where texts overlapped with spatial objects, the list-based approach was often preferred, highlighting the role of spatial organization in effective information presentation.
Other non-data elements other than annotation could also potentially augment human perceptual experience. Minimizing ink, as originally proposed by Tufte [135], may not be the ideal approach for designing graphics. For example, adding a background grid generally improves accuracy in quantitative tasks [50]. The minimalist approach contrasts with other strategies, such as incorporating chart-junks [104, 57] or subtle contextual cues (e. g., [148]), which can augment human visualization perceptual experience [37]).
V.3 Considerations of Image-based Method to Capture Task-Free Perceptual Experiences
In this work, we showed a quantitative approach that combines human subjective readings and objective metric measurement to study the mechanism underlying complex perceptual experiences. We aimed to create a representative and relatively large dataset with a broad coverage, while also minimizing human expert curation. Here, we have demonstrated how current data can be used to study “unobserved” complexity through the modeling of limited metrics and quantitative method. While big data has historically been used to address challenges in other disciplines, we can now leverage our own (relatively) big data to tackle computational problems by working with visualization-specific inquiries—the answer is always to serve and better understand human perception—while quantifying important underlying mechanisms.
Because this approach relies on data-driven significance and interpretive assessments across many factors, it helps identify a set of potential interpretable predictors. The goal, however, is not to claim that our quantitative metrics fully capture the richness of human perceptual experience or task-specific uses. Rather, this method measures judgments at scale, by linking perceived and objective factors. The use of large, real-world image datasets enables us to capture a wide range of design variations, establishing a robust baseline for understanding cognitive processing in visualization, akin to the memorability experiment [11]. When carefully chosen, these metrics can effectively reflect perceptual judgments [72]. As a result, models built on these features can offer a principled foundation for extracting high-level perceptual experiences from visualization inputs, and for measuring how these experiences change in response to visual design variations.
@@ V.3 Considerations of Image-based Method to Capture Task-Free Perceptual Experiences
We recommend that progress towards establishing mechanisms underlying VC through quantitative analysis should address three key challenges: first, the development of common and comprehensive image databases to support the assessment of perceptual experiences [113]; second, the design of experiments that directly capture high-level, perceptually driven experiences, such as meaningfulness, understandability, etc; third, the collection and validation of objective metrics capable of quantifying those perceptual experiences.
@@
Our results revealed that human high-level understanding is supported not only by design-level stimuli but also by efficient object-level representations. The implication is to extend the current visualization representation units—whether derived from visual design or cognitive perception—by showing how they are formed by segmenting an image into understandable chucks. This reduction process is generally efficient and this understanding process is likely driven by visual complexity, which balances representational cost and memory load. By correlating complexity scores with model performance on visualization understanding tasks, one could also automatically flag overly complex figures for downstream processing, extending the impact of our work into the era of AI-driven data analysis.
@@ V.4 Limitations and Future Work
Despite the broader range of metrics and assessments included in this study, limitations remain. Experiments like ours are often limited by the datasets used and the tasks given. For example, we could also have validated our outcomes using a small dataset through item response theory, similar to approaches used in literacy testing [105], for diagnostic efficiency. Here we prioritize explainability based on large data and objective metrics. In addition, our study used images free of interaction or animation, viewed by the general public rather than domain experts. One can use our method to incorporate different modalities when engaging in new datasets. Finally, A most exciting direction for future research would be to extend feature congestion modeling [117] to account for stimulus-specific complexity beyond continuous color-texture stimulus. In principle, metric-based modeling offers the potential for greater precision in capturing a large set of variables of how complexity is perceived.
@@ VI Conclusion
Our empirical study systematically examines the relationship between human perception and objective image quality metrics. The VisComplexity2K dataset contains visualization stimulus types, human-rated VC scores, and 12 metric measurement scores. Our analyses reveal that the number of elements, whether defined by low-level features like edges and points or high-level boundaries formed by color and text, is a key factor in perceived visual complexity and shapes whether a visualization is perceived as scene-like.
@@
Ultimately, such an approach could provide a principled basis for interpreting high-level perceptual experiences across diverse visualization inputs, thereby enhancing the scientific and design potential of complexity quantification. Many aspects of visualization experience and metric behavior extend beyond what is observable in controlled lab-based experiments. This underscores the need to study real-world visualizations used in everyday contexts. To that end, the methodology used in this study, collecting subjective human ratings through large-scale crowdsourcing, offers a scalable and ecologically valid way to capture holistic human perception in real-world settings. The complementary insights gained from combining data-driven modeling and experimental methods have the potential to drive transformative advances - an opportunity that the visualization research community is well positioned to embrace.
@@
We, as authors, state that Figure 1, Figure 5, Figure 6, Figure 7, right pane, and Figure 8, right pane, as well as Table 2 and Table 1 are under our own copyright with the permission to be used here. We have also made them available under the Creative Commons Attribution 4.0 International (Γ Δ CC BY 4.0) license and share them at osf.io/bdet6. All remaining images in the paper are © IEEE, with permission to be used here.
@@ Acknowledgements.
The work was partially supported by the “Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2120/1 – 390831618” through a Visiting Professorship Position, made to Prof. Dr. Jian Chen, via the Cluster of Excellent IntCDC, University of Stuttgart, Germany. The crowdsourcing experiment was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 251654672 – TRR 161, with The Ohio State University IRB approval ID: 2022B0363. We also wish to thank Dr. Ruth Rosenholtz at MIT for sharing the Matlab code for computing O.FC, and colleagues at the Visual Attention Lab at Harvard University and reviewers for their comments.References
@@
- [3] Y. A. Andrienko, N. Brilliantov, and J. Kurths. Complexity of two-dimensional patterns. Eur. Phys. J. B, 15:539–546, 2000. doi: 10/c6bzm7
- [4] A. Ashokkumar and J. W. Pennebaker. Social media conversations reveal large psychological shifts caused by covid-19’s onset across us cities. Science advances, 7(39):eabg7843, 2021.
@@
- [8] G. Barua, T. Pingel, and T. Lim. Urban thermal map design considerations: color, shading, and resolution. Cat & Geo Info Sci, 51(4):549–564, 2024. doi: doi.org/pd8k
@@
- [10] M. Borkin, Z. Bylinskii, G. Krzysztof, N. Kim, A. Oliva, and H. Pfister. Massachusetts (massive) visualization dataset (massvis).
@@
- [26] S. Corchs, F. Gasparini, and R. Schettini. No reference image quality classification for jpeg-distorted images. Dgt Sigl Proc, 30, 2014. doi: 10/f54jpv
- [27] S. E. Corchs, G. Ciocca, E. Bricolo, and F. Gasparini. Predicting complexity perception of real world images. PLoS one, 11(6), 2016. doi: 10/gbpnrm
@@
- [30] D. DeCarlo, A. Finkelstein, S. Rusinkiewicz, and A. Santella. Suggestive contours for conveying shape. In Seminal Graphics Papers: Pushing the Boundaries, Volume 2, pp. 401–408. ACM, 2023.
- [31] K. Dinkla, M. A. Westenberg, and J. J. van Wijk. Compressed Adjacency Matrices: Untangling gene regulatory networks. IEEE TVCG, 18(12), 2012. doi: 10/f4fwnz
@@
- [33] D. C. Donderi. Visual complexity: a review. Psychological bulletin, 132(1):73, 2006. doi: 10/frf92v
- [34] Y. Du, C. Li, R. Guo, X. Yin, W. Liu, J. Zhou, Y. Bai, Z. Yu, Y. Yang, Q. Dang, et al. PP-OCR: A Practical Ultra Lightweight OCR System. ArXiv, 2020. doi: 10/nvxk
@@
- [36] A. Farhadi, I. Endres, D. Hoiem, and D. Forsyth. Describing objects by their attributes. In Proc. CVPR, pp. 1778–1785, 2009. doi: 10/cwdbn7
@@
- [38] C. Fernandez-Lozano, A. Carballal, P. Machado, A. Santos, and J. Romero. Visual complexity modelling based on image features fusion of multiple kernels. PeerJ, 7, 2019. doi: 10/mzzt
@@
- [40] R. M. Gray. Entropy and information theory. Sci. Bus. Media, 2011. doi: 10/gnms9f
- [41] M. R. Greene. The information content of scene categories. In Psychol. Learn. Motiv., vol. 70. Elsevier, 2019. doi: 10/nvxg
@@
- [46] D. Hasler and S. E. Suesstrunk. Measuring colorfulness in natural images. In Hum Vis Electron Imaging VIII, vol. 5007, pp. 87–95. SPIE, 2003. doi: 10/bj98jc
@@
- [50] J. Heer and M. Stone. Color naming models for color selection, image editing and palette design. In Proc the ACM SIGCHI, pp. 1007–1016, 2012. doi: 10/gfz5q4
@@
- [56] D. E. Huber and C. G. Healey. Visualizing data with motion. In IEEE VIS, 2005., pp. 527–534, 2005. doi: 10/fjpgkx
@@
- [63] P. Isola, J. Xiao, A. Torralba, and A. Oliva. What makes an image memorable? In Proc. CVPR, pp. 145–152, 2011. doi: 10/cznr72
@@
- [69] B. Kerr. Thread arcs: An email thread visualization. In IEEE Symposium on Information Visualization 2003 (IEEE Cat. No. 03TH8714), pp. 211–218. IEEE, 2003. doi: 10/bsb89f
@@
- [72] M. A. Kramer, M. N. Hebart, C. I. Baker, and W. A. Bainbridge. The features underlying the memorability of objects. Science advances, 9(17), 2023. doi: 10/gr72gs
@@
- [74] C. Kyle-Davidson, O. Solis, S. Robinson, R. T. W. Tan, and K. K. Evans. Scene complexity and the detail trace of human long-term visual memory. Vision Research, 227:108525, 2025. doi: 10/pfj3
@@
- [76] H. Lam, E. Bertini, P. Isenberg, C. Plaisant, and S. Carpendale. Empirical Studies in Information Visualization: Seven Scenarios. IEEE TVCG, 18, 2012. doi: 10/drrh6j
- [77] P. Q. Le, A. M. Iliyasu, J. Garcia, F. Dong, and K. Hirota. Representing visual complexity of images using a 3D feature space based on structure, noise, and diversity. J Adv Comp Intell, 16(5), 2012. doi: 10/mzz6
@@
- [86] M. L. Mack and A. Oliva. The perceptual dimensions of visual simplicity. J. Vis, 4(719), 2004. doi: 10/frr5nb
- [87] I. Mario, M. Chacon, D. Alma, and S. Corral. Image complexity measure: a human criterion free approach. In Annu. Meet. North Am. Fuzzy Inf. Process. Soc., pp. 241–246, 2005. doi: 10/bdqq23
@@
- [89] T. Mehmood, K. H. Liland, L. Snipen, and S. Sæbø. A review of variable selection methods in partial least squares regression. Chemometrics and intelligent laboratory systems, 118:62–69, 2012.
@@
- [91] S. Miksch, H. Leitte, and M. Chen. Knowledge-Assisted Visualization and Guidance. Found. Data Vis., pp. 61–85, 2020. doi: 10/nvxn
- [92] A. Miniukovich and A. De Angeli. Quantification of interface visual complexity. In Proc. Inter. Working Conf. on Adv. Visual Interfaces, pp. 153–160, 2014. doi: 10/gm7pmd
@@
- [94] T. Minka, R. Cleven, and Y. Zaykov. Trueskill 2: An improved bayesian skill rating system. Tech Rep, 2018.
@@
- [97] F. Nagle and N. Lavie. Predicting human complexity perception of real-world scenes. R Soc Open Sci, 7(5):191487, 2020. doi: 10/gq6f3m
@@
- [100] A. Oliva and A. Torralba. Modeling the Shape of the Scene: A Holistic Representation of the Spatial Envelope. Int J Comput Vis, 42:145–175, 2001. doi: 10/dm4d3b
- [101] A. Oliva and A. Torralba. Building the gist of a scene: The role of global image features in recognition. Prog. Brain Res., 155:23–36, 2006. doi: 10/b2jzn8
- [102] M. A. Otaduy, N. Jain, A. Sud, and M. C. Lin. Haptic display of interaction between textured models. In ACM SIGGRAPH 2005 Courses, pp. 133–es, 2005. doi: 10/b95rph
- [103] S. Palan and C. Schitter. Prolific.ac—A subject pool for online experiments. J. Behav. Exp. Finance, 17:22–27, 2018. doi: 10/gftct9
- [104] L. Panavas, A. E. Worth, T. Crnovrsanin, T. Sathyamurthi, S. Cordes, M. A. Borkin, and C. Dunne. Juvenile Graphical Perception: A Comparison between Children and Adults. In ACM Hum-Comput Interact, article no. 138, 14 pages, pp. 1–14, 2022. doi: 10/nvzd

### v3

@@ IV Comparison to Previous Works
IV.1 Network-based Visual Complexity: Compared to Purchase [107]
@@
To replicate and understand this extended large dataset, and compare results with the Purchase et al. study on network aesthetics, we curated a subset of our dataset containing network-style visualizations (primarily node-link diagrams, as shown in Apdx. Figure 24). To better match the original study design, we manually (1) filtered for direct node-link diagram (such that the matrix views were removed); (2) Removed all text and annotations using automated masking followed by a manual cleanup to isolate pure graphical elements given that these texts may activate the edge detector. We refer to this cleaned subset as HP-Trees and Networks. Using our 11 remaining metrics (excluding O.TiR), our PLS regression achieved an of 0.58 indicating a strong fit—substantially higher than the 0.25 reported in Purchase et al. [107]. Notably, O.ED (), O.FP (), O.MeC (), and O.CF () were key contributors to VC in this subset (Figure 7). We would credit the results to the addition of O.MeC and O.FP, which emerge as main contributors to perceived VC of networks.
IV.2 Information-Theoretic Measure of Color and Texture on Clutter Effect: Compared to Rosenholtz et al. [120]
Rosenholtz et al. designed this clutter-based metric to capture color and texture distributions in a natural scene. Our dataset contained 311 ‘grid and matrix’ or ‘heatmaps’ [10], of which 185 are spatial, continuous, where spatial location ‘is given’ [95], making them more closely aligned with Rosenholtz’s definition of structured scenes. The rest are discrete, with pixel locations artificially generated—where structured grids and matrix introduce corners and edges. This contrast is evident in our dataset in these two sets of heatmaps (Figure 8). We also removed all text before modeling these two sets of visualization images to avoid edges and corners confounded by the presence of text.
@@
In summary, the continuity-specific results largely mirror the overall findings in terms of dominant contributing metrics, with some variation based on the stimulus type the metric models. What makes this particularly intriguing is that O.FC is highly effective in explaining complexity for continuous heatmap images but had limited explanatory power for artificially generated heatmaps. Given that the coefficients are higher by individual visualization types and continuity, we recommend that type categorization for VC should take into account continuity.
@@ V General Discussion
This section discusses theoretical and design implications of our results. A central problem for the science of visualization is related to perceptual experience as it influences people’ understanding about data. One of such high-level perceptual experiences is visual complexity. Our present work contributes to understanding the alignment between human reading and metric measurement.
@@ V.1 Comparisons to Natural Scene Understanding in terms of Item-Count and Feature Congestion
Multiscale structures emerge in our study. Number of elments operates on object-associated O.MeC and O.TiR, and low-level features such as edges O.ED and corners O.FP, contributing significantly to VC across overall dataset. This result indicates that human perception does not operate on a single scale [153]. This high-level perception seems to align to Turing’s distinct hypothesis: patterned structures can self-organize through local competitive interactions [138]. As a result, perception can be fast and minimize the need for generic instruction to perform pattern findings. Turing described that people’s perceptual computing in this competing mechanism appeared to be more robust to noise, where global structural emergence exhibited topological robustness of insensitivity to variations and noises.
In visualization contexts, global structures formed by color and pattern may support visual Gestalt principles, as captured by the two object-based metrics: O.TiR (related to text) and O.MeC (related to color and object boundaries). These size-scale representations showed that low-level elements also emerge at the level of higher-order textures and objects. This is particularly evident in heatmaps where spatial location is given and aligns with the descriptor attribute O.FC, as demonstrated in our experiments (Figure 8). In this context, O.FC represents a meaningful step forward in quantifying human perceptual experiences by treating certain visualization types (here color and texture) as statistical, scene-based representations.
This result may also indirectly support two perspectives. First, while Borkin et al. [10] treated grid and matrix as a single visualization type to characterize perceptual experiment (memorability), our findings may support that continuity can be considered as a distinguishing factor when we define type. Second, the prolific investigations of information-theoretic measures in data visualizations [22, 65, 63, 157] may contribute to humans’ perceptual experiences—particularly when the underlying information bits are well defined. In our case, the color/texture dimensions in O.FC can be interpreted as encoding structured visual patterns. This result further confirms that continuous color/texture patterns in visualizations is understood differently from discrete grid and matrix.
@@ V.2 Graphic or Text? The Influence of Semantic Information on Perceived Visual Complexity
Our findings suggest that the semantic information conveyed by text in visualizations impacts perceived VC differently than graphic elements. We conjectured that explanatory text may reduce perceived VC by enhancing understandability, e. g., Hearst et al. [47] also highlighted the significance of text, considering it as equally important as the visualizations themselves. We found a positive effect of text on VC to a certain extent. In general, explanatory text reduces perceived VC due to its contribution to understandability. However, when the amount of text becomes excessive, it may increase complexity, shifting from aiding comprehension to overwhelming the visual features. Thus using titles and short annotations would be better than long paragraphs.
@@
Analyzing variations in text layout may reveal new ways to understand perceived VC. Previous eye-tracking behavior data for capturing memorability suggested that people can quickly distinguish text and titles. These text elements always occupied separate areas on the screen, making them easy for the human visual system to decode. In previous studies on three-dimensional scenes, where text overlapped with spatial objects, the list-based approach was often preferred [20], highlighting the role of spatial organization in effective information presentation. Other non-data elements beyond annotations may also augment human perceptual experience, although their interaction with graphical elements can be complex. For example, adding a background grid [133], incorporating chart-junk [55], or using subtle contextual cues [150] have all been shown to augment human visualization perceptual experience [36]. The question of how much text is sufficient, and how non-data elements, graphical features, and annotations should appear may need to be personalized to align with the observers’ reasoning goals [104] and the narrative structure of the visualization.
V.3 Considerations of Data-Driven Image-based Data Collection
@@
The use of large, real-world image datasets enabled us to capture a wide range of metric representations. When carefully chosen, a set of metrics can effectively reflect perceptual judgments in other large-scale studies [70]. We thus recommend that progress towards establishing mechanisms underlying VC through quantitative analysis could use large-data collection to address three key challenges: first, the development of common and comprehensive image databases to support the assessment of perceptual experiences [113]; second, the design of experiments that directly capture high-level, perceptually driven experiences, such as meaningfulness, understandability, etc; third, the collection and validation of objective metrics capable of explaining those perceptual experiences that would eventually inform design.
@@
Our results revealed that human high-level understanding is supported not only by design-level stimuli (e. g., point and line) but also by efficient object-level representations (e. g., color and text). The implication is to extend the current visualization representation units—whether derived from visual design or cognitive perception—by showing how they are formed by segmenting an image into understandable subsets. Each of these units is mapped to a corresponding data item, enabling fixation allocation and interaction while preserving essential information for the human observer. Finally, one could also automatically flag overly complex figures for downstream processing, extending the impact of our work into the era of AI-driven data analysis.
@@ V.4 Limitations and Future Work
Despite the broader range of metrics and assessments included in this study, limitations remain. Experiments like ours are often limited by the datasets used. For example, one could also validate our outcomes using a small dataset through item response theory for diagnostic efficiency, similar to approaches used in literacy testing [103]. Here we prioritize explainability based on large data and objective metrics. In addition, our study used images free of interaction or animation, viewed by the general public rather than domain experts. Also, our evaluation is task-free. One can use our method to incorporate different modalities. Finally, an exciting direction for future research would be to extend feature congestion modeling [118] to account for stimulus-specific complexity beyond continuous color-texture stimulus. In principle, metric-based modeling offers the potential for greater precision in capturing a large set of variables of how complexity is perceived.
@@ VI Conclusion
Our empirical study systematically examines the relationship between human perception and objective image quality metrics. The go.osu.edu/viscomplexitydata contains visualization stimulus types, human-rated VC scores, and 12 metric measurement scores. Our analyses reveal that the number of elements, whether defined by low-level features like edges and points or high-level boundaries formed by color and text, is a key factor in perceived visual complexity and shapes whether a visualization is perceived as scene-like.
@@
Ultimately, such an approach could provide a principled basis for interpreting high-level perceptual experiences across diverse visualization inputs, thereby enhancing the scientific and design potential of complexity quantification. Many aspects of visualization experience and metric behavior extend beyond what is observable in controlled lab-based experiments. This underscores the need to study real-world visualizations used in everyday contexts. To that end, the methodology used in this study, collecting subjective human ratings through large-scale crowdsourcing, offers a scalable way to capture holistic human perception in real-world settings. The complementary insights gained from combining data-driven modeling and experimental methods have the potential to drive transformative advances—an opportunity that the visualization research community is well-positioned to embrace.
@@
We, as authors, state that Figures 1, 5, 6, 7, right pane, and 8, right pane, as well as Table 2 and Table 1 are under our own copyright with the permission to be used here. We have also made them available under the Creative Commons Attribution 4.0 International (Γ Δ CC BY 4.0) license and share them at osf.io/bdet6. All remaining images in the paper are © IEEE, with permission to be used here.
@@ Acknowledgements.
The work was partially supported by the “Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2120/1 – 390831618” through a Visiting Professorship Position, made to Prof. Dr. Jian Chen, via the Cluster of Excellent IntCDC, University of Stuttgart, Germany. The crowdsourcing experiment was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 251654672 – TRR 161, with The Ohio State University IRB approval ID: 2022B0363. R.S. Laramee was supported, in part, by funding from the EPSRC (EPSRC UKRI157, APP17227). We also wish to thank Dr. Ruth Rosenholtz at MIT for sharing the Matlab code for computing Feature Congestion, and colleagues at the Visual Attention Lab at Harvard University, as well as reviewers for their comments.References
@@
- [3] Y. A. Andrienko, N. Brilliantov, and J. Kurths. Complexity of two-dimensional patterns. Eur Phys J B, 15:539–546, 2000. doi: 10/c6bzm7
- [4] A. Ashokkumar and J. W. Pennebaker. Social media conversations reveal large psychological shifts caused by COVID-19’s onset across US cities. Science advances, 7(39):eabg7843, 2021. doi: 10/gm6vps
@@
- [8] G. Barua, T. Pingel, and T. Lim. Urban thermal map design considerations: color, shading, and resolution. Cat & Geo Info Sci, 51(4):549–564, 2024. doi: 10/pd8k
@@
- [20] J. Chen, P. S. Pyla, and D. A. Bowman. Testbed evaluation of navigation and text display techniques in an information-rich virtual environment. In IEEE VR, pp. 181–289, 2004. doi: 10/dch9x2
@@
- [26] S. Corchs, F. Gasparini, and R. Schettini. No reference image quality classification for JPEG-distorted images. Dgt Sigl Proc, 30:86–100, 2014. doi: 10/f54jpv
- [27] S. E. Corchs, G. Ciocca, E. Bricolo, and F. Gasparini. Predicting complexity perception of real world images. PLoS one, 11(6):e0157986, 2016. doi: 10/gbpnrm
@@
- [30] D. DeCarlo, A. Finkelstein, S. Rusinkiewicz, and A. Santella. Suggestive contours for conveying shape. In Seminal Graphics Papers: Pushing the Boundaries, vol. 2, pp. 401–408. ACM, 2023. doi: 10/bpk44q
- [31] K. Dinkla, M. A. Westenberg, and J. J. van Wijk. Compressed Adjacency Matrices: Untangling gene regulatory networks. IEEE TVCG, 18(12):2457–2466, 2012. doi: 10/f4fwnz
@@
- [33] D. C. Donderi. Visual complexity: A review. Psycho Bulletin, 132(1):73, 2006. doi: 10/frf92v
@@
- [35] A. Farhadi, I. Endres, D. Hoiem, and D. Forsyth. Describing objects by their attributes. In Proc CVPR, pp. 1778–1785, 2009. doi: 10/cwdbn7
@@
- [37] C. Fernandez-Lozano, A. Carballal, P. Machado, A. Santos, and J. Romero. Visual complexity modelling based on image features fusion of multiple kernels. PeerJ, 7:e7075, 2019. doi: 10/mzzt
@@
- [39] M. R. Greene. The information content of scene categories. In Psychol Learn Motiv, vol. 70, pp. 161–194. Elsevier, 2019. doi: 10/nvxg
@@
- [44] D. Hasler and S. E. Suesstrunk. Measuring colorfulness in natural images. In Hum Vis Elec Imaging VIII, vol. 5007, pp. 87–95. SPIE, 2003. doi: 10/bj98jc
@@
- [48] J. Heer and M. Stone. Color naming models for color selection, image editing and palette design. In ACM SIGCHI, pp. 1007–1016, 2012. doi: 10/gfz5q4
@@
- [54] D. E. Huber and C. G. Healey. Visualizing data with motion. In IEEE Vis, pp. 527–534, 2005. doi: 10/fjpgkx
@@
- [61] P. Isola, J. Xiao, A. Torralba, and A. Oliva. What makes an image memorable? In Proc CVPR, pp. 145–152, 2011. doi: 10/cznr72
@@
- [67] B. Kerr. Thread Arcs: An email thread visualization. In IEEE InfoVis, pp. 211–218, 2003. doi: 10/bsb89f
@@
- [70] M. A. Kramer, M. N. Hebart, C. I. Baker, and W. A. Bainbridge. The features underlying the memorability of objects. Nature Sci Adv, 9(17):eadd2981, 2023. doi: 10/gr72gs
@@
- [72] C. Kyle-Davidson, O. Solis, S. Robinson, R. T. W. Tan, and K. K. Evans. Scene complexity and the detail trace of human long-term visual memory. Vision Res, 227:108525, 2025. doi: 10/pfj3
@@
- [74] H. Lam, E. Bertini, P. Isenberg, C. Plaisant, and S. Carpendale. Empirical studies in information visualization: Seven scenarios. IEEE TVCG, 18(9):1520–1536, 2012. doi: 10/drrh6j
- [75] P. Q. Le, A. M. Iliyasu, J. Garcia, F. Dong, and K. Hirota. Representing visual complexity of images using a 3D feature space based on structure, noise, and diversity. J Adv Comp Intell, 16(5):631–640, 2012. doi: 10/mzz6
@@
- [77] C. Li, W. Liu, R. Guo, X. Yin, K. Jiang, Y. Du, Y. Du, L. Zhu, B. Lai, X. Hu, D. Yu, and M. Yanjun. PP-OCRv3: More attempts for the improvement of ultra lightweight OCR system. ArXiv, 2022. doi: doi:10/gtjnh4
@@
- [85] M. L. Mack and A. Oliva. The perceptual dimensions of visual simplicity. J Vision, 4(719):10–1167, 2004. doi: 10/frr5nb
- [86] I. Mario, M. Chacon, D. Alma, and S. Corral. Image Complexity Measure: a human criterion free approach. In Annu Meet North Am Fuzzy Inf Process Soc, pp. 241–246, 2005. doi: 10/bdqq23
@@
- [88] T. Mehmood, K. H. Liland, L. Snipen, and S. Sæbø. A review of variable selection methods in partial least squares regression. Chemometrics and intelligent laboratory systems, 118:62–69, 2012. doi: 10/f4hx5x
@@
- [90] S. Miksch, H. Leitte, and M. Chen. Knowledge-assisted visualization and guidance. Found Data Vis, pp. 61–85, 2020. doi: 10/nvxn
- [91] A. Miniukovich and A. De Angeli. Quantification of interface visual complexity. In Proceedings of the 2014 international working conference on advanced visual interfaces, pp. 153–160, 2014. doi: 10/gm7pmd
@@
- [93] T. Minka, R. Cleven, and Y. Zaykov. TrueSkill 2: An improved bayesian skill rating system. Technical report, Microsoft Research, 2018.
@@
- [96] F. Nagle and N. Lavie. Predicting human complexity perception of real-world scenes. Royal Soc Open Sci, 7(5):191487, 2020. doi: 10/gq6f3m
@@
- [99] A. Oliva and A. Torralba. Modeling the Shape of the Scene: A holistic representation of the spatial envelope. Int J Comput Vision, 42:145–175, 2001. doi: 10/dm4d3b
- [100] A. Oliva and A. Torralba. Building the gist of a scene: The role of global image features in recognition. Prog Brain Res, 155:23–36, 2006. doi: 10/b2jzn8
- [101] M. A. Otaduy, N. Jain, A. Sud, and M. C. Lin. Haptic display of interaction between textured models. In ACM SIGGRAPH Courses, pp. 133–es, 2005. doi: 10/b95rph
- [102] S. Palan and C. Schitter. Prolific.ac—a subject pool for online experiments. J Behav Exp Finance, 17:22–27, 2018. doi: 10/gftct9